7.The family of all the subsets of any set S is called
a.the power set of S
b.the null set of S
c.the identity set of S
d.the cardinality set of S

8.Let E = {2, 4, 6, ...}. What is the compliment of the set E?
a.odd numbers
b.prime numbers
c.even numbers
d.rational numbers
1
Expert's answer
2019-09-10T10:57:16-0400

The family of subsets of any set "S" is called the power set of "S", so the answer to question 7 is a).

Assuming that the universe is the set of natural numbers "\mathbb{N} = {1, 2, 3, ...}" , by definition, the complement of E is "E^c = \mathbb{N} \setminus E = 1, 3, 5, ..." - the set of odd numbers, so the answer to question 8 is a).